OpenStudy (anonymous):

Consider the equations y=1/√x and y=x−2. The approximate solution of this system of equations is ____

OpenStudy (welshfella):

replace y by 1/ sqrt x 1/ sqrt x = x - 2 square both sides 1 /x = (x-2)^2 can you continue/

OpenStudy (welshfella):

1 is one of the roots if we replace x by 1 in the above equation we get 1/1 = (-1)^2 1 = 1

OpenStudy (welshfella):

but that does'nt fit the original equation so its not 1